# Distributed Evolution Strategies with Multi-Level Learning for Large-Scale Black-Box Optimization

### v1

@@ III-C Spatiotemporal Global Step-Size Adaptation (STA)
In order to exploit both the spatio and temporal (non-local) information on-the-fly, our meta-framework combines the less-known Meta-ES strategy with the well-known CSA strategy (or its recent population-based or success-based variants for LSO)888In this paper, we do not modify the CSA-style strategy for all inner-ESs. Instead, we focus on the global step-size adaptation at the outer-ES level, which is crucial for DES.. Specifically, three parallel design strategies are used at the outer-ES to balance adaptation speed and meta-population diversity[143]: 1) when the elitist strategy is used for the distribution mean update, the same elitist strategy is also used to update the global step-size of the corresponding inner-ES to obtain the relatively stable evolution process; 2) given a predefined proportion (e.g., 1/5), some inner-ESs mutate the weighted multi-recombined global step-size as , where , implicitly based on the strong causality assumption 999For simplicity, in this paper we follow the common suggestion from Meta-ES and set and , respectively (note that leads to unbiasedness in the logarithmic scale).; 3) otherwise, the global step-size will be uniformly sampling from a reasonable search range101010In practice, a reasonable search range seems to be easier to set than a reasonable value., in order to maintain diversity and reduce the risk of getting trapped into local minima.
@@ III-D Collective Learning of CMA on Structured Populations
The most prominent feature of CMA-ES appears to be its adaptive encoding (i.e., invariance against affine transformation) ability, especially on non-separable, ill-conditioned problems. As a general-purpose black-box optimizer, we expect DES to keep this powerful feature as much as possible. In order to be communication-efficient, however, we need to properly compress the standard covariance matrix to fit for the distributed shared memory, but which may destroy the highly desirable invariance property. In this paper, we choose to use one of its large-scale variants (i.e., LM-CMA) as the basic computing unit on each CPU core, in order to reduce the communication cost after each isolation time .
@@
For properly approximating CMA under distributed computing, our meta-framework employs two adaptive distributed strategies for structure and distribution learning, respectively. First, as the structure learning often operates at a relatively slow-changing scale and controls the richness of distribution model, we implicitly adapt the total number of reconstructed evolution paths via the elitist strategy, in order to obtain a reliable learning progress at the outer-ES level. In other worlds, we save a certain elitist ratio111111It is set to for consistency and simplicity in this paper. as some (but not all) parallel inner-ESs for the next isolation time and for the remaining parallel inner-ESs we sample uniformly in a reasonable setting range121212In practice, the setting of the search range of depends on the available memory in the used distributed computing platform, which is easy to obtain., in order to maintain the diversity of structure learning at the outer-ES level. Note that for each inner-ES, its reconstructed structure is always fixed within each isolation time.
@@ III-E A Meta-Framework for DES
Here we combine all the aforementioned design choices into our distributed meta-framework, as presented in Algorithm 1131313Rinnooy Kan and Timmer [144] from the mathematical programming community considered a multi-level method for stochastic optimization in the context of single linkage. However, our approach is orthogonal to their method.. To implement our meta-framework, we select one state-of-the-art clustering computing software called ray[145] as the key engine of distributed computing. As compared with other existing distributed computing systems such as MPI [146], P2P [147, 148], MapReduce [149, 15], Spark [150], BlockChain [151], ray provides a flexible programming interface for Python and a powerful distributed scheduling strategy to cater to modern challenging AI applications such as population-based training [152, 153], AutoML [154], and open-ended learning/evolution [155]. Owing to the intrinsic complexity of distributed algorithms, we provide an open-source Python implementation for our proposed meta-framework available at https://github.com/Evolutionary-Intelligence/M-DES, in order to ensure repeatability and benchmarking[156].
For simplicity and ease to analyze, our meta-framework uses the generational (rather than steady-state) population update strategy at the outer-ES level. Although typically the steady-state method could maximize the parallelism level especially for heterogeneous environments, the asynchronous manner makes distributed black-box optimizers often difficult to debug. In this paper, we consider only the generational population update manner, since it makes the updates and communications of search/mutation distributions easier to understand and analyze under distributing computing.
@@ IV Large-Scale Numerical Experiments
To study the benefits (and costs) of our proposed meta-framework (simply named as DLMCMA here), we conduct numerical experiments on a set of large-scale benchmarking functions with memory-expensive fitness evaluations, arguably reflecting many challenging data-driven optimization problems. To ensure repeatability and promote benchmarking [157], all involved experimental data and Python code are openly available at our companion website https://github.com/Evolutionary-Intelligence/M-DES.
@@ IV-A Experimental Settings
Test Functions: We choose a set of 13 commonly used test functions, as shown in Table I for their mathematical formula (see e.g., COCO/BBOB [22] or NeverGrad [158] for implementations). These functions can be roughly classified as two families (i.e., unimodal and multimodal functions) to compare local and global search abilities, respectively. For benchmarking large-scale BBO, the dimensions of all the test functions are set to 2000. We also use the standard angle-preserving (i.e., rotation) transformation [17] (rather than [24]) and random shift to generate non-separability and avoid the origin as the global optimum, respectively. This involved matrix-vector multiplication operator results in the memory-expensive fitness evaluation, arguably one significant feature of many real-world data-driven optimization problems. In order to speedup parallel fitness evaluations, we use the simple yet efficient shared memory trick for our distributed optimizer.
@@
Benchmarking Optimizers: To benchmark the advantages and disadvantages of different approaches, we select a total of 61 black-box optimizers from different families (i.e., Evolution Strategies - ES, Natural Evolution Strategies - NES, Estimation of Distribution Algorithms - EDA, Cross-Entropy Method - CEM, Differential Evolution - DE, Particle Swarm Optimizer - PSO, Cooperative Coevolution - CC, Simulated Annealing - SA, Genetic Algorithms - GA, Evolutionary Programming - EP, Pattern Search - PS, and Random Search - RS) implemented in a recently well-designed Python library called PyPop7. Due to the page limit, here we only use their abbreviated (rather full) names to avoid legend confusion when plotting the convergence figures. For details about these selected optimizers, please refer to this online website pypop.rtfd.io and references therein. Overall, these optimizers are grouped into two classes for plotting convergence curves: ES-based optimizers (excluding NESs) and others (including NESs), as shown in Fig. (1, 3) and Fig. (2, 4), respectively.
Computing Environments: We run all experiments on two private clustering computing platforms with different CPU-core scales for unimodal and multimodal functions, respectively. We set the parallel number of inner-ESs () to 380 and 540 (CPU cores) for our DLMCMA on unimodal and multimodal functions, respectively. On multimodal functions, more parallel inner-ESs often bring more population diversity. Although the optimal setting of is problem-dependent, here we focus on the scalability w.r.t. CPU cores and do not fine-tune it for each function. We set the isolation time to 150 seconds, based on our experiments (it is not necessarily an optimal value).
Owing to the time-consuming experiment process for LSO, We run each optimizer 10 and 4 times for each unimodal and multimodal function, respectively. The total CPU core runtime to be needed in our experiments is estimated up to 18600 hours, that is, 775 days = hours.
@@ IV-B Comparing Local Search Abilities
On the sphere function, arguably one of the simplest test cases for continuous optimization, it is highly expected that the optimizer could obtain a fast convergence. Most ES-based optimizers could obtain a satisfactory (not necessarily optimal) performance except some CMA-ES variants with quadratic complexity (Fig. 1a). For quadratic-complexity CMA-ES variants (e.g., DD-CMA [119], MA-ES [122], and C-CMA-ES [124]), the overall runtime is dominated heavily by the CMA mechanism rather than the function evaluation time, therefore resulting in a much slow adaptation speed.
@@
There is one predominated search direction needed to explore for the cigar function. This means that a low-rank learning strategy (e.g., R1-ES [116]) is typically enough to capture the main direction via adaptation. Owing to the extra cost brought from distributed computing, our meta-framework (DLMCMA) obtains a slightly slow convergence speed as shown in Fig. 1b (reflecting limitations of parallelism [159]). However, it could approximate the low-rank learning ability well, given that the initial number of reconstructed evolution paths does not match the optimal setting.
@@
For both functions discus and ellipsoid, there exist multiple promising search directions (see their relatively even eigenvalue distributions). Therefore, a much richer reconstruction model is preferred for CMA. On the discus function (Fig. 1c), our DLMCMA could show the 3x runtime speed w.r.t. the second ranked optimizer (i.e., MM-ES [121]). On the ellipsoid function, our DLMCMA nearly always shows the best convergence speed during evolution (Fig. 1d), because its collective learning strategy maintains the better diversity of reconstructed evolution paths via utilizing the distributed computing resource. Interestingly, similar observations could also be found in another two challenging functions (i.e., differentpowers and schwefel12) with multiple search directions (as shown in Fig. 1e and 1f, respectively).
@@
For both schwefel221 and step functions, there are a large number of plateaus for the high-dimensional cases, which result in a rugged fitness landscape. For ESs, a key challenge is to properly adapt the global step-size to pass these plateaus141414In other worlds, this needs to find an appropriate evolution window, popularized by Rechenberg (one of the evolutionary computation pioneers).. Luckily, our meta-framework can tackle this challenge well and can obtain the best convergence rate on both of them (Fig. 1g and 1h), since our STA strategy could keep the diversity of the global step-size well while avoiding to diverge it (with the help of the elitist strategy)151515The punctuated-equilibria-style convergence is dated back to at least [160], depending on the used viewpoint (as pointed out by Schwefel, one of the evolutionary computation pioneers)..
@@
As we can clearly see from Fig. 3, our DLMCMA could always obtain the best convergence rate when compared with all other algorithm families except that on schwefel221 R1-NES shows a very similar performance. Overall, our DLMCMA can obtain the best or competitive performance on these unimodal function, empirically validating the benefits from multilevel distributed learning.
@@ IV-C Comparing Global Search Abilities
For minimizing the ackley function, it seems to be looking for a needle in the haystack. However, there is a global landscape structure to be available, which can be used to accelerate the global convergence rate of the optimizer (if well-designed). We find that many large-scale variants of CMA-ES utilize this property, even given a small population setting. Our DLMCMA can approximate this global structure well and therefore it obtains the best convergence rate after bypassing all (shallow) local optima (Fig. 3a and 4a).
On the classic rastrigin function, there exist a large number of relatively deep local minima, which can hinder the optimization process. To escape from these local optima, the simple yet efficient restart [161] strategy from CMA-ES will increase the number of offspring after each restart. Clearly, our DLMCMA obtains much better results among all optimizers, with the help of multiple restarts (Fig. 3b and 4b).
Our proposed DLMCMA obtains the second ranking only after UMDA [162] on the michalewicz function (Fig. 3c and 4c), which seems to have a relatively weak global structure. The default population size of UMDA is relatively large (200) while that of each (local) LM-CMA used in our DLMCMA is very small (19) by default. Despite this difference, our DLMCMA still can drive the parallel evolution process over structured populations to approach the best after 3 hours.
On the multimodal function salomon, our D-LM-CMA finds the best solution much faster than all other optimizers (Fig. 3d and 4d). However, the restart strategy cannot help to find a better solution, which may means that this found solution is near a deep local optimum. On another multimodal function scaledrastrigin, our DLMCMA ranks the third, only after R1-NES and SNES [91] (Fig. 3e and 4e). We notice that the parallel evolution process stagnates even at the early stage, which means that we need a better restart strategy for this function. We leave it for future work.
@@ V Conclusion
In this paper, we propose a multilevel learning-based meta-framework to parallelize one large-scale variant of CMA-ES called LM-CMA, significantly extending our previous conference paper [163]. Within this meta-framework, four main design choices are made to control distribution mean update, global step-size adaptation, and CMA reconstruction for scalability and effectiveness. A large number of comparative experiments show the benefits (and costs) of our proposed meta-framework.

### v4

@@ III-C Spatiotemporal Global Step-Size Adaptation (STA)
In order to exploit both the spatio and temporal (non-local) information on-the-fly, our meta-framework combines the less-known Meta-ES strategy with the well-known CSA strategy (or its recent population-based or success-based variants for LSO)888In this paper, we do not modify the CSA-style strategy for all inner-ESs. Instead, we focus on the global step-size adaptation at the outer-ES level, which is crucial for DES.. Specifically, three parallel design strategies are used at the outer-ES to balance adaptation speed and meta-population diversity[141]: 1) when the elitist strategy is used for the distribution mean update, the same elitist strategy is also used to update the global step-size of the corresponding inner-ES to obtain the relatively stable evolution process; 2) given a predefined proportion (e.g., 1/5), some inner-ESs mutate the weighted multi-recombined global step-size as , where , implicitly based on the strong causality assumption 999For simplicity, in this paper we follow the common suggestion from Meta-ES and set and , respectively (note that leads to unbiasedness in the logarithmic scale).; 3) otherwise, the global step-size will be uniformly sampled from a reasonable search range101010In practice, a reasonable search range seems to be easier to set than a reasonable value., in order to maintain diversity and reduce the risk of getting trapped into local minima.
@@ III-D Collective Learning of CMA on Structured Populations
The most prominent feature of CMA-ES appears to be its adaptive encoding (i.e., invariance against affine transformation) ability, especially on non-separable, ill-conditioned problems. As a general-purpose black-box optimizer, we expect DES to keep this powerful feature as much as possible. In order to be communication-efficient, however, we need to properly compress the standard covariance matrix to fit the distributed shared memory; but this may destroy the highly desirable invariance property. In this paper, we choose to use one of its large-scale variants (i.e., LM-CMA) as the basic computing unit on each CPU core, in order to reduce the communication cost after each isolation time .
@@
For properly approximating CMA under distributed computing, our meta-framework employs two adaptive distributed strategies for structure and distribution learning, respectively. First, as the structure learning often operates at a relatively slow-changing scale and controls the richness of distribution model, we implicitly adapt the total number of reconstructed evolution paths via the elitist strategy, in order to obtain a reliable learning progress at the outer-ES level. In other words, we save a certain elitist ratio111111It is set to for consistency and simplicity in this paper. as some (but not all) parallel inner-ESs for the next isolation time and for the remaining parallel inner-ESs we sample uniformly in a reasonable setting range121212In practice, the setting of the search range of depends on the available memory in the used distributed computing platform, which is easy to obtain., in order to maintain the diversity of structure learning at the outer-ES level. Note that for each inner-ES, its reconstructed structure is always fixed within each isolation time.
@@ III-E A Meta-Framework for DES
Here we combine all the aforementioned design choices into our distributed meta-framework, as presented in Algorithm 1131313Rinnooy Kan and Timmer [142] from the mathematical programming community considered a multi-level method for stochastic optimization in the context of single linkage. However, our approach is orthogonal to their method.. To implement our meta-framework, we select one state-of-the-art clustering computing software called ray[143] as the key engine of distributed computing141414https://github.com/Evolutionary-Intelligence/M-DES/blob/main/README.md#troubleshooting-tips. As compared with other existing distributed computing systems such as MPI [144], P2P [145, 146], MapReduce [147, 15], Spark [148], BlockChain [149], ray provides a flexible programming interface for Python and a powerful distributed scheduling strategy to cater to modern challenging AI applications such as population-based training [150, 151], AutoML [152], and open-ended learning/evolution [153]. Owing to the intrinsic complexity of distributed algorithms, we provide an open-source Python implementation for our proposed meta-framework available at https://github.com/Evolutionary-Intelligence/M-DES, in order to ensure repeatability and benchmarking[154].
For simplicity and ease to analyze, our meta-framework uses the generational (rather than steady-state) population update strategy at the outer-ES level. Although typically the steady-state method could maximize the parallelism level especially for heterogeneous environments, the asynchronous manner makes distributed black-box optimizers difficult to debug. In this paper, we consider only the generational population update manner, since it makes the updates and communications of search/mutation distributions easier to understand and analyze under distributing computing.
@@ IV Large-Scale Numerical Experiments
To study the benefits (and costs) of our proposed meta-framework (simply named as DLMCMA here), we conduct numerical experiments on a set of large-scale benchmarking functions with memory-expensive fitness evaluations, arguably reflecting many challenging data-driven optimization problems. To ensure repeatability151515https://github.com/Evolutionary-Intelligence/M-DES/blob/main/README.md#step-by-step-instructions and promote benchmarking [155], a set of involved experimental data and Python code are openly available at our companion website https://github.com/Evolutionary-Intelligence/M-DES.
@@ IV-A Experimental Settings
Test Functions: We choose a set of 13 commonly used test functions, as shown in Table II for their mathematical formula (see e.g., COCO/BBOB [22] or NeverGrad [156] for implementations). These functions can be roughly classified to two families (i.e., unimodal and multimodal functions) to compare local and global search abilities, respectively. For benchmarking large-scale BBO, the dimensions of all the test functions are set to 2000. We also use the standard angle-preserving (i.e., rotation) transformation [17] (rather than [24]) and random shift to generate non-separability and avoid the origin as the global optimum, respectively. This involved matrix-vector multiplication operator results in the memory-expensive fitness evaluation, arguably one significant feature of many real-world data-driven optimization problems. In order to speedup parallel fitness evaluations, we use the simple yet efficient shared memory trick for our distributed optimizer.
@@
Benchmarking Optimizers: To benchmark the advantages and disadvantages of different approaches, we select a total of 61 black-box optimizers from different families (i.e., Evolution Strategies - ES, Natural Evolution Strategies - NES, Estimation of Distribution Algorithms - EDA, Cross-Entropy Method - CEM, Differential Evolution - DE, Particle Swarm Optimizer - PSO, Cooperative Coevolution - CC, Simulated Annealing - SA, Genetic Algorithms - GA, Evolutionary Programming - EP, Pattern Search - PS, and Random Search - RS) implemented in a recently well-designed Python library called PyPop7. Due to the page limit, here we only use their abbreviated (rather than full) names to avoid legend confusion when plotting the convergence figures. For implementation details about all these optimizers and their hyper-parameter settings, please refer to this online website pypop.rtfd.io and references therein. Overall, these optimizers are grouped into two classes for plotting convergence curves: ES-based optimizers and others, as shown in Fig. (2, 4) and Fig. (3, 5), respectively.
Computing Environments: We set the parallel number of inner-ESs () to 380 and 540 (CPU cores) for our DLMCMA on unimodal and multimodal functions, respectively. On multimodal functions, more parallel inner-ESs often bring larger population diversity. Although the optimal setting of is problem-dependent, we do not fine-tune it for each function. We empirically set the isolation time to 150 seconds on all functions though this is not necessarily an optimal value for each function.
Owing to the time-consuming experiment process for LSO, We run each optimizer 10 and 4 times on each unimodal and multimodal function, respectively. The total CPU single-core runtime needed in our experiments is estimated up to 18600 hours, that is, 775 days = hours.
@@ IV-B Comparing Local Search Abilities
On the sphere function, arguably one of the simplest test cases for continuous optimization, it is highly expected that the optimizer could obtain a fast rate of convergence. Most ES-based optimizers could obtain a satisfactory (not necessarily optimal) performance except some CMA-ES variants with quadratic complexity (Fig. 2). For quadratic-complexity CMA-ES variants (e.g., DD-CMA [121], MA-ES [60], and C-CMA-ES [61]), the overall runtime is dominated heavily by the CMA mechanism rather than the function evaluation time, therefore resulting in a much lower adaptation speed.
@@
There is one predominated search direction needed to be explored for the cigar function. This means that a low-rank learning strategy (e.g., R1-ES [57]) is typically enough to capture the main direction via adaptation. Owing to the extra cost brought from distributed computing, our meta-framework (DLMCMA) obtains a slightly slow convergence speed as shown in Fig. 2 (reflecting limitations of parallelism [157]). However, it could approximate the low-rank learning ability well, given that the initial number of reconstructed evolution paths does not match the optimal setting.
@@
For both functions discus and ellipsoid, there exist multiple promising search directions (see their relatively even eigenvalue distributions). Therefore, a much richer reconstruction model is preferred for CMA. On the discus function (Fig. 2), our DLMCMA could show the 3x runtime speedup w.r.t. the second ranked optimizer (i.e., MM-ES [56]). On the ellipsoid function, our DLMCMA nearly always shows the best convergence speed during evolution (Fig. 2), because its collective learning strategy maintains the better diversity of reconstructed evolution paths via utilizing the distributed computing resource. Interestingly, similar observations could also be found in another two challenging functions (i.e., differentpowers and schwefel12) with multiple search directions (Fig. 2).
@@
For both schwefel221 and step functions, there are a large number of plateaus in high-dimensional cases, which result in a rugged fitness landscape. For ESs, a key challenge is to properly adapt the global step-size to pass these plateaus161616In other words, this needs to find an appropriate evolution window, popularized by Rechenberg (one of the evolutionary computation pioneers).. Luckily, our meta-framework can tackle this challenge well and can achieve the best convergence rate on both of them (Fig. 2), since our STA strategy could keep the diversity of the global step-size well while avoiding to diverge it (with the help of the elitist strategy)171717The punctuated-equilibria-style convergence is dated back to at least [158], depending on the used viewpoint (as pointed out by Schwefel, one of the evolutionary computation pioneers)..
@@
As we can clearly see from Fig. 3, our DLMCMA could always obtain the best convergence rate when compared with all other algorithm families except that on schwefel221 R1-NES shows a very similar performance. Overall, our DLMCMA achieves the best or competitive performance on these unimodal functions, validating the benefits of multilevel distributed learning empirically.
@@ IV-C Comparing Global Search Abilities
For minimizing the ackley function, it seems to be looking for “a needle in the haystack”. However, there is a global landscape structure to be available, which can be used to accelerate the global convergence rate of the optimizer (if well-utilized). We find that many large-scale variants of CMA-ES utilize this property, even under a small population setting. Our DLMCMA can also approximate this global structure well, therefore achieving the best convergence rate after bypassing all (shallow) local optima (Fig. 4 and 5).
On the classic rastrigin function, there exist a large number of relatively deep local minima, which can hinder the optimization process. To escape from these local optima, the simple yet efficient restart [159] strategy from CMA-ES will increase the number of offspring after each restart. Clearly, our DLMCMA obtains much better results among all optimizers, with the help of multiple restarts (Fig. 4 and 5).
Our proposed DLMCMA obtains the second ranking only after UMDA [160] on the michalewicz function (Fig. 4 and 5), which seems to have a relatively weak global structure. The default population size of UMDA is relatively large (200) while that of each (local) LM-CMA used in our DLMCMA is small by default. Despite this difference, our DLMCMA still can drive the parallel evolution process over structured populations to approach the best after 3 hours.
On the multimodal function salomon, our DLMCMA finds the best solution much faster than all other optimizers (Fig. 4 and 5). However, the restart strategy cannot help to find a better solution, which may indicate that this found solution is near a deep local optimum. On another multimodal function scaledrastrigin, our DLMCMA ranks the third, only after R1-NES and SNES [99] (Fig. 4 and 5). We notice that the parallel evolution process stagnates even at the early stage, which means that we need a better restart strategy for this function. We leave it for future work.
@@
IV-D Overhead Analysis of Memory Communications
For our distributed algorithm, a set of evolution paths, distributed over different nodes, are needed to reconstruct the covariance matrix. If this set is too large, it can lead to expensive communication overheads. Otherwise, if this set is too small, it may damage the model richness of the reconstructed covariance matrix. As a result, a trade-off between communication overheads and model richness should be made properly. We can calculate181818github.com/Evolutionary-Intelligence/M-DES/blob/main/figures/plot_overhead_of_memory_communications.py the amount of memory communications over the network under different settings of number of evolution paths (e.g., 100, 500, 1000, and 2000), as shown in Fig. 6 (on a 2000-dimensional fitness function). Clearly, a high compress ratio (e.g., 100/2000) can significantly reduce the amount of memory communications especially when the level of parallelism is high, while achieving the best performance on most cases (see Figs. 2, 3, 4, and 5 for empirical demonstrations).
IV-E Trade-off Analysis of Performance
In this subsection, we discuss performance trade-offs of our distributed algorithm with one case study called black-box classification from data science (with a non-convex tanh loss function). For this loss function, a business dataset from the popular UCI Machine Learning Repository191919https://doi.org/10.24432/C51G7P is employed, leading to an 857-dimensional fitness function (to be minimized).
As we can see from Fig. 7 and Fig. 8a, given four different levels of parallelism (that is, 1, 100, 200, and 300), the higher the level of parallelism, the faster (more) the convergence (number of function evaluations). To keep the memory amount to be communicated at a reasonable level, we choose a low number (i.e., 25 in our experiments) as the maximum of evolution paths, as shown in Fig. 8b. If the full-ranked covariance matrix was reconstructed (i.e., the number of evolution paths is set to 857), the memory amount to be communicated over the network would increase by x times, which could result in a significant communication overhead for our distributed algorithm.
@@ V Conclusion
In this paper, we propose a multilevel learning-based meta-framework to parallelize one large-scale variant of CMA-ES called LM-CMA, significantly extending our previous conference paper [161]. Within this meta-framework, four main design choices are made to control distribution mean update, global step-size adaptation, and CMA reconstruction for effectiveness and efficiency. A large number of comparative experiments show the benefits (and costs) of our proposed meta-framework.